OpenStudy (anonymous):

some one can solve this...lim ((arctanx/x)/lnx) x→∞ some one can solve this...lim ((arctanx/x)/lnx) x→∞ @Mathematics

OpenStudy (across):

\[\lim_{x\to\infty}\frac{\frac{\arctan x}{x}}{\ln x}.\]This?

OpenStudy (anonymous):

yes

OpenStudy (across):

Well, this can be rewritten as\[\lim_{x\to\infty}\frac{\arctan x}{x\ln x}\implies\lim_{x\to\infty}\frac{\arctan x}{\ln x^x}.\]The limit of the arctangent as this approaches infinity is pi/2, and the limit of ln x^x as this approaches infinity is infinity. Therefore, the limit of the whole function is 0.

OpenStudy (anonymous):

is that your answer

OpenStudy (anonymous):

i think the answer is pi/2

OpenStudy (across):

\[\lim_{x\to\infty}\arctan x=\frac{\pi}{2}.\]\[\lim_{x\to\infty}x\ln x=\infty.\]

OpenStudy (anonymous):

you have to use L`hopital regel,,,,and you will get pi/2

OpenStudy (across):

Ich habe denn keine Ahnung, was deine ursprüngliche Frage ist.

OpenStudy (anonymous):

Oh....i cant speak duch

OpenStudy (across):

Und warum benutzt du die Regel von L'Hospital, wenn der Limes des Zählers existiert?

OpenStudy (across):

Why are you using L'Hôpital's rule when the limit of the numerator exists?

OpenStudy (anonymous):

let me ask you whre u from,,?

OpenStudy (across):

I'm German.

OpenStudy (anonymous):

iam from norway,,

OpenStudy (anonymous):

wenn Sie sehen, der Zähler des Ausdrucks, so dass Sie unendliche x 0 haben

OpenStudy (across):

Ich konnte das nicht verstehen. Aber man kann nur diese Regel benutzen, wenn man das hat: \[\lim_{x\to c}f(x)=\lim_{x\to c}g(x)=0,\]oder\[\lim_{x\to c}f(x)=\lim_{x\to c}g(x)=\pm\infty,\]und\[\lim_{x\to c}\frac{f'(x)}{g'(x)}\]existiert.

OpenStudy (anonymous):

look the quation again,,,and how it writens,,,

OpenStudy (anonymous):

i mean before you take x to the numinator

OpenStudy (anonymous):

prove that limx→∞ (arctanx/x ) = ∞. solve this one